Example. A stretched string of length 50cm is set oscillating by displacing its mid distance of 2cm from its rest point a equation-=- where c21, to determine the resulting motion u(x, t). 2 2 iKnx) 25 50 x(cm)
Answer #1

Here,length of the string is 50cm after stretching it and given wave equation is

(d2u/dx2)=(1/c2)(d2u/dt2). Where d is a partial differential operator

We can also write this equation as

X"t=(1/c2)T"x.

x"/x=t"/t=c2.

And substitute the given value of c2 in the above differential equation

This can be written as

x"/x=t"/t=1.

And by separation of variables ODE equations become

x"-x=0. and t"-t=0.

We can simplify above equations as

(D2-1)x=0 and (D2-1)t=0.

As D=+-1 for both x and t,General solution will become

u(x,t)=(c1ex+c2e-x)(c3et+c4e-t).

Initial Conditions from the graph

u(0,t)=0 and u(x,0)=0

u(50,t)=(2/25)(e25(2)-e-25(-2))=(2/25)(e50-e-50).It is triangular wave.

u(0,t)=c1+c2=0 and u(x,0)=c3+c4=0.

u(50,t)=c1e50+c2e-50=(2/25)(e50-e-50).and we got c1=-c2.

After substituting above conditions

We will get c1=-(2/25) and c2=(2/25).

u(2,x)=f(x)=(2/25)(2-25)=(2/25)(-23)=(c3e2+c4e-2) and c3=-c4

and c3=(2/25)(-23)=-46/25.

Substitute above values and we get general solution as

u(x,t)=2/25(e-50-e50)(-46/25)(e-2-e2).